# D-LGP: Dynamic Logic-Geometric Program for Combined Task and Motion Planning

### v1

@@ Abstract
Many real-world sequential manipulation tasks involve a combination of discrete symbolic search and continuous motion planning, collectively known as combined task and motion planning (TAMP). However, prevailing methods often struggle with the computational burden and intricate combinatorial challenges stemming from the multitude of action skeletons. To address this, we propose Dynamic Logic-Geometric Program (D-LGP), a novel approach integrating Dynamic Tree Search and global optimization for efficient hybrid planning. Through empirical evaluation on three benchmarks, we demonstrate the efficacy of our approach, showcasing superior performance in comparison to state-of-the-art techniques. We validate our approach through simulation and demonstrate its capability for online replanning under uncertainty and external disturbances in the real world.
@@ I Introduction
Robots are increasingly playing a vital role in real-world scenarios, particularly with long-horizon manipulation tasks. These tasks can usually be decoupled with various subtasks, each associated with continuous trajectories. This refers to a problem called combined Task and Motion Planning (TAMP), where high-level discrete task planning and low-level continuous motion planning are tightly interconnected. The effectiveness of the motion planner heavily relies on the task skeleton, and, simultaneously, the task planner should be aware of how its outputs are evaluated by the motion planner to avoid spending excessive time on bad solutions. Due to this inherent coupling, the entire system, encompassing all objects and robots, must be considered holistically. This can be viewed as a high-dimensional underactuated kinematic system [1].
In this high-dimensional space, there are several submanifolds corresponding to different object configurations, with some low-dimensional intersected subspaces shared among them. Therefore, as highlighted by Garrett et al. [1], TAMP can be more accurately described as a hybrid discrete-continuous search problem. It aims to find a sequence of discrete mode types (e.g., which objects to pick and place) along with the corresponding continuous motion paths within each mode and the configuration parameters in the intersection space.
@@
Figure 1: System setup where the robot constructs a tower sequentially. The objective is to stack the blocks at a specified target point in a predefined order. If a block is out of reach (indicated by the red curve), the robot must determine how to use a tool to pull the block into the reachable region before performing the pick-and-place operation.
@@
While discrete and continuous planning have been studied separately for a long time in the fields of AI and robotics, TAMP tries to combine aspects of both symbolic AI planning and geometric robot motion planning. Many advances have been made to address this integration challenge, typically involving a search-based strategy to identify task skeletons and sampling-based or optimization-based motion planning methods to evaluate them. The interplay between the task and motion domains results in significant high-dimensional combinatorial complexity, particularly as the number of objects and the length of trajectory horizon increase. Unlike continuous planning, where dimensionality can be easily reduced thanks to the interpolation nature, such as using basis functions [2], there is no direct numerical relationship between adjacent symbolic variables in TAMP. This means that even a small change in a high-level task skeleton can lead to a significant difference in the continuous domain. Consequently, this combinatorial complexity of TAMP leads to unavoidable high-dimensional solution space, posing significant challenges for high-level task planners to find a feasible candidate from a huge amount of choices. For example, popular discrete planning methods like Monte Carlo Tree Search (MCTS) suffer a lot from the sparse solution space, despite efforts to introduce multiple bounds [3] or approximated cost-to-go functions [4]. Low-level motion planning, typically sampling-based or optimization-based, is used to evaluate the quality of the selected action skeleton. Nevertheless, these evaluations often lack guarantees of reliability. Sampling-based techniques only provide probabilistic completeness under infinite time, which is impractical for TAMP to be solved in real time. They yield feasible but non-optimal solutions, while optimality is important for real-world robot manipulation, such as time and energy efficiency. While optimization-based approaches are geared towards optimality, they primarily rely on nonlinear programming (NLP) [5]. The resulting local optima can sometimes be satisfactory, but they may lack the confidence to generate the best solution and provide certain feedback about feasibility to high-level task planners. In summary, the intricate high-dimensional combinatorial complexities lead to inefficient task sampling, exacerbated by the poor performance of low-level motion planning. These challenges collectively classify TAMP as an NP-hard problem, rendering the quest for a feasible solution time-consuming. This hinders the online replanning for robot manipulation, which is important for robots to interact with the real world.
@@
In this work, we address these challenges by introducing a method that uses backpropagation for high-level task skeleton reasoning and global optimization for low-level motion planning. The task planner employs the principles of Dynamic Programming [6], where the action skeleton is derived through backpropagation from the target configuration. This backward mechanism also eliminates limitations on horizon lengths. Consequently, we refer to it as Dynamic Tree Search (DTS). In terms of low-level motion planning, we consider non-convex motion planning as a combination of many convex subspaces, allowing us to formulate the problem as mixed-integer convex optimization and solve it globally using modern solvers.
@@
3) Given the fast computation of D-LGP, we formulate it in a receding-horizon planning fashion that can handle uncertainty and external disturbances in real-world scenarios.
@@ II Related Work
Task and motion planning is initiated with single-flow approaches like aSyMov [7] and SMAP [8], that start from task planning to motion planning. These methods require a complete description of constraints but lack interplay between the task layer and the motion layer. To address this limitation, more TAMP methods have emerged, allowing low-level motion planners to validate the feasibility of task skeletons. Examples include PDDLStream [9] and FFRob [10]. These methods incorporate geometric constraints as logical predicates, demanding extensive human effort in logic design. The success of these methods relies heavily on the accuracy of the underlying models, which may not always capture the nuances of the real-world environment.
Recognizing that the feasibility of task sequences is validated by the resulting motion planning, TAMP approaches have evolved to allow the task planner to receive and react to feedback from the motion planner during the planning process. This iterative refinement approach is exemplified in works like [11, 12]. Failures in motion planning are captured as logical predicates, aiding in the creation of new skeletal structures to avoid geometric issues. However, the iterative nature of this approach results in multiple rounds of communication and coordination between the high-level and low-level planners, which can be time-consuming. The incorporation of human expertise to discretize the continuous motion space for symbolic expression introduces apprehensions regarding the completeness of planning algorithms.
@@
Among the methods mentioned above, [9, 10, 12] primarily fall into the category of sampling-based TAMP, utilizing existing sampling planners like RRTs [13] to assess the geometric feasibility of selected action skeletons. Sampling-based methods theoretically achieve probabilistic completeness given unlimited time. However, TAMP typically requires fast motion planning to guide task planners. Moreover, sampling-based methods prioritize feasibility over optimality, while this is vital for efficient robot manipulation in terms of time and energy.
@@
This leads to the emergence of optimization-based TAMP, with Logic-Geometric Programming (LGP) [5] being a prominent framework. LGP employs Monte-Carlo Tree Search (MCTS) in the STRIPS domain to identify action skeleton candidates and subsequently solves nonlinear trajectory optimization problems with path and mode switching constraints. However, LGPs often exhibit slow runtimes due to the number of calls to the NLP solver. In cases where the search space is sparse, most candidate action skeletons prove infeasible, resulting in a protracted and unsatisfactory exploration process. Several variants have been proposed to mitigate this. For instance, Toussaint et al. introduced multi-bound tree search (MBTS) [3], which solves a relaxed NLP problem that serves as a lower bound for the original problem and a necessary condition for feasibility. Building upon this concept, more variants have been put forward, each employing different bounds and heuristics to prune infeasible branches or approximate cost-to-go functions of MCTS. Examples include stability bounds [14], prefix conflicts [15], and action-specific heuristics [4]. However, these bounds and heuristics are often domain-specific and depend on human knowledge. Even when well-designed, these search strategies may still struggle with the sparsity of the solution space, leading to the exploration of irrelevant action skeletons and unavoidable time consumption. Unless the cost-to-go functions can be perfectly approximated, which is usually not the case, these challenges persist. Overall, MCTS-based strategies are highly effective for most tasks when given sufficient time. However, they are usually too general to capture the causal structure of certain TAMP problems, particularly those with precisely known target configurations. Such problems are prevalent in the field of robot manipulation, including tabletop rearrangement and assembly tasks.
@@
Our approach falls under the LGP category and aims to achieve efficient and globally optimal TAMP for robot manipulation tasks in which the target configuration is known. We employ backward search to fully uncover the causal structure of long-horizon tasks, enhancing the efficiency of task planning. The feasibility checking adopts the multi-bound concept from [3] for fast computation. We formulate motion planning problems as mixed-integer convex optimization, efficiently solvable with modern solvers, such as Gurobi [16], to obtain global optimal solutions or determine infeasibility. This hierarchical structure results in an efficient and globally optimal LGP framework, providing fast and reliable feedback from the geometric level to the symbolic level, ultimately enabling our approach to find feasible action skeletons and globally optimal solutions.
@@ III-A Target-centric LGP
The full path of LGP is divided into $K$ phases of a fixed duration $T$, with $s_{k}$ representing the logical states at phase $k$. Instead of checking feasibility and computing the cost of full path after finding the full skeleton candidate, we propose to use backward propagation from the target configuration. This requires a new target-centric LGP formulation:
@@
$h_{\text{switch}}(\bm{x}(t_{k})|a_{k},s_{k-1})=0$
$g_{\text{switch}}(\bm{x}(t_{k})\,|\,a_{k},s_{k-1})\leq 0$
@@
where $h_{\text{path}}$ and $g_{\text{path}}$ represent the path constraints under each phase, and $h_{\text{switch}}$, $g_{\text{switch}}$ denote the transition conditions between two phases with the action operator $a_{k}$. $g_{k}$ is the target to be achieved at phase $k$. $s_{k}\in\mathcal{G}(g_{k})$ and $h_{\text{goal}}$ describes the goal specification logically and geometrically. $c(.)$ and $c_{T}(.)$ are the path cost and terminal cost to penalize control effort (e.g., energy, force) and the final kinematic performance. The logical state transition from $s_{k-1}$ to $s_{k}$ is determined by $succ(.,.)$ as a function of $s_{k-1}$ and the discrete action $a_{k}\in\mathcal{A}(s_{k-1})$. $\mathcal{T}^{\dagger}(g_{k})$ identifies the essential subgoals to attain $g_{k}$, which is described in detail in Alg. 1. The formulation (1) adopts the concepts and variables from the standard LGP formulation [5], while allowing dynamic programming for backpropagation from the target configuration $g_{K}$. This mechanism can be viewed as spatial reasoning over the known configuration structure of the environments.
@@ III-B Mixed-integer Motion Planning
The goal of tabletop rearrangement is to manipulate blocks to achieve a desired configuration. This task involves a set of objects denoted as $\mathcal{O}$, each with poses represented as $P_{n}^{t}=(x_{n}^{t},y_{n}^{t},\theta_{n}^{t})\in SE(2)$ at time $t$ for object $n$. In essence, the problem entails determining the optimal placement $\bm{u}_{t}=[x_{j}^{t}\;y_{j}^{t}]^{\scriptscriptstyle\top}$ for a selected object $j$, which can be thought of as a form of obstacle avoidance. This task is notably challenging due to the non-convex nature of the feasible points beyond closed and bounded squares. Rather than attempting to solve the problem within this non-convex set, we adopt a strategy of decomposition into multiple convex sets. Subsequently, the constraints for each object $n\in 1,\dots,N$ can be succinctly expressed as combinations of four halfspaces, denoted by $i\in\{1,2,3,4\}$:
$\bm{C_{i}}:\quad$ $[a_{i}\quad b_{i}]\;\bm{R}(\theta_{n}^{t})^{\scriptscriptstyle\top}\;\bm{u}_{t}\geq c_{i},$ (2)
where $\bm{R}(\theta_{i}^{t})$ is a rotation matrix of the block frame and the global frame:
$R(\theta_{i}^{t})=\begin{bmatrix}\text{cos}(\theta_{i}^{t})&-\text{sin}(\theta_{i}^{t})\\ \text{sin}(\theta_{i}^{t})&\text{cos}(\theta_{i}^{t})\end{bmatrix},$
and $a_{1}=1,b_{1}=0,c_{1}=l/2+x_{n}^{t}$, $a_{2}=-1,b_{2}=0,c_{2}=l/2-x_{n}^{t}$, $a_{3}=0,b_{3}=1,c_{3}=l/2+y_{n}^{t}$, $a_{4}=0,b_{4}=-1,c_{4}=l/2-y_{n}^{t}$. $l$ is the size of the square block.
As a consequence, the non-intersection constraints among these four halfspaces can be viewed as a disjunctive constraint $\lor_{i=1}^{4}\bm{C}_{i}$, which can be further represented by combining constraints along with the introduction of additional binary variables, as discussed in [17].
The comprehensive formulation can be converted into a Mixed-Integer Quadratic Program (MIQP) characterized by a quadratic objective and linear constraints. This transformation incorporates binary variables denoted as $z_{i}$ to determine the specific halfspace within which the placement should reside. In the case of each halfspace, the constraints can be expressed linearly. To articulate the if-else condition related to binary variables, we can employ a conventional big-M formulation, similar to [11], to represent (2) as
$\bm{C_{n}}:\quad$ $[a_{i}\quad b_{i}]\;\bm{R}(\theta_{n}^{t})^{\scriptscriptstyle\top}\;\bm{u}_{t}\geq c_{i}-M(1-z_{i})$ (3)
$\sum z_{i}=1,\quad z_{i}\in\{0,1\},\quad\forall i\in\{1,2,3,4\}$
Given action skeleton $a_{k}$, the motion planning problem in (1) can be described as
@@
$\forall t\in[(k-1)T,kT]:\bm{u}(t)\in\cup_{n=1}^{N}\{\bm{C_{n}}\}.$

### v3

# D-LGP: Dynamic Logic-Geometric Program for Reactive Task and Motion Planning

@@ Abstract
Many real-world sequential manipulation tasks involve a combination of discrete symbolic search and continuous motion planning, collectively known as combined task and motion planning (TAMP). However, prevailing methods often struggle with the computational burden and intricate combinatorial challenges, limiting their applications for online replanning in the real world. To address this, we propose Dynamic Logic-Geometric Program (D-LGP), a novel approach integrating Dynamic Tree Search and global optimization for efficient hybrid planning. Through empirical evaluation on three benchmarks, we demonstrate the efficacy of our approach, showcasing superior performance in comparison to state-of-the-art techniques. We validate our approach through simulation and demonstrate its reactive capability to cope with online uncertainty and external disturbances in the real world.
@@ I Introduction
Robots are increasingly playing a vital role in real-world scenarios, particularly with long-horizon manipulation tasks. These tasks can usually be decoupled with various subtasks, each associated with continuous trajectories. This refers to a problem called combined Task and Motion Planning (TAMP) [1], where high-level discrete task planning and low-level continuous motion planning are tightly interconnected. The effectiveness of the motion planner heavily relies on the task skeleton, and, simultaneously, the task planner should be aware of how its outputs are evaluated by the motion planner to avoid spending excessive time on bad solutions. Due to this inherent coupling, the entire system, encompassing all objects and robots, must be considered holistically.
@@
Figure 1: System setup where the robot constructs a tower sequentially. The objective is to stack the blocks at a specified target point in a predefined order. If a block is out of reach (indicated by the red curve), the robot must determine how to use a tool to pull the block into the reachable region before continuing the pick-and-place operation.
@@
Many advances have been made to address this integration challenge in TAMP, typically involving a search-based strategy to identify task skeletons and sampling-based or optimization-based motion planning methods to evaluate them. The interplay between the task and motion domains results in significant high-dimensional combinatorial complexity, particularly as the number of objects and the length of trajectory horizon increase. Unlike continuous planning, where dimensionality can be easily reduced thanks to the interpolation nature, such as using basis functions [2], there is no direct numerical relationship between adjacent symbolic variables in TAMP. This means that even a small change in the high-level task skeleton can lead to a significant difference in the continuous domain. Consequently, this combinatorial complexity of TAMP leads to unavoidable high-dimensional solution space, posing significant challenges for the high-level task planner to find a feasible skeleton from a huge amount of choices. For example, popular discrete planning methods like Monte Carlo Tree Search (MCTS) suffer a lot from the sparse solution space, despite efforts to introduce multiple bounds [3] or approximated cost-to-go functions [4]. Low-level motion planning, typically sampling-based or optimization-based, is used to evaluate the quality of the selected action skeleton. Nevertheless, these evaluations often lack guarantees of reliability. Sampling-based techniques are probabilistically complete given infinite time, which is impractical for TAMP to be solved in real time. They yield feasible but non-optimal solutions, while optimality is important for real-world manipulation, such as time and energy efficiency. Optimization-based approaches are geared towards optimality, they primarily rely on nonlinear programming (NLP) [5]. The resulting local optima can sometimes be satisfactory, but they may lack the confidence to provide certain feedback about feasibility to high-level task planners. In summary, the intricate high-dimensional combinatorial complexities lead to inefficient task sampling, exacerbated by the poor performance of low-level motion planning. These challenges collectively classify TAMP as an NP-hard problem, rendering the quest for a feasible solution time-consuming. This hinders the online replanning for robot manipulation, which is important for robots to interact with the real world reactively.
@@
In this work, we address these challenges by introducing a method that uses backpropagation for high-level action skeleton reasoning and global optimization for low-level motion planning. The task planner employs the optimal principles of Dynamic Programming [6], where the action skeleton is derived through backpropagation from the target configuration. This backward mechanism also eliminates limitations on horizon lengths. Consequently, we refer to it as Dynamic Tree Search (DTS). In terms of low-level motion planning, we consider non-convex motion planning as a combination of many convex subspaces, allowing us to formulate the problem as mixed-integer convex optimization and solve it globally using modern solvers.
@@
3) Given the fast computation of D-LGP, we formulate it in closed-loop fashion that is reactive to uncertainty and external disturbances in real-world scenarios.
@@ II Related Work
Task and motion planning is initiated with single-flow approaches like aSyMov [7] and SMAP [8], that start from task planning to motion planning. These methods require a complete description of constraints but lack interplay between the task layer and the motion layer. To address this limitation, more TAMP methods have emerged, allowing low-level motion planners to validate the feasibility of task skeletons. Examples include PDDLStream [9] and FFRob [10]. These methods incorporate geometric constraints as logical predicates, demanding extensive human efforts in logic design. The incorporation of human expertise introduces apprehensions regarding the algorithm completeness.
@@
Among the methods mentioned above, [9, 10] primarily fall into the category of sampling-based TAMP, utilizing existing sampling planners like RRTs [11] to assess the geometric feasibility of selected action skeletons. Sampling-based methods theoretically achieve probabilistic completeness given unlimited time [12]. However, TAMP typically requires fast motion planning to inform task planners. Moreover, sampling-based methods prioritize feasibility over optimality, while the latter is vital for efficient robot manipulation in terms of time and energy.
@@
This leads to the emergence of optimization-based TAMP, with Logic-Geometric Programming (LGP) [5] being a prominent framework. LGP employs Monte-Carlo Tree Search (MCTS) in the STRIPS domain to identify action skeleton candidates and subsequently solves nonlinear trajectory optimization problems with path and mode switching constraints. However, LGPs often exhibit slow runtimes due to the number of calls to the NLP solver. In cases where the search space is sparse, most selected action skeletons are usually infeasible, resulting in a protracted and unsatisfactory exploration process. Several variants have been proposed to mitigate this. For instance, Toussaint et al. introduced multi-bound tree search (MBTS) [3], which solves a relaxed NLP problem that serves as a lower bound for the original problem and a necessary condition for feasibility. Building upon this concept, more variants have been put forward, each employing different bounds and heuristics to prune infeasible branches or approximate cost-to-go functions of MCTS. Examples include stability bounds [13], prefix conflicts [14], and action-specific heuristics [4]. However, these bounds and heuristics are often domain-specific and depend on human knowledge. Even when well-designed, these search strategies may still struggle with the sparsity of the solution space, leading to the exploration of irrelevant action skeletons and unavoidable time consumption. Unless the cost-to-go functions can be perfectly approximated, which is usually not the case, these challenges persist. Overall, MCTS-based strategies are highly effective for most tasks when given sufficient time. However, they are usually too general to capture the causal structure of certain TAMP problems, particularly those with precisely known target configurations. Such problems are prevalent in the field of robot manipulation, including tabletop rearrangement and assembly tasks.
@@
Our approach falls under the LGP category and aims to achieve efficient and globally optimal TAMP for robot manipulation tasks in which the target configuration is known. We employ backward search to fully uncover the causal structure of long-horizon tasks, enhancing the efficiency of task planning. The feasibility checking adopts the multi-bound concept from [3] for fast computation. We formulate motion planning problems as mixed-integer convex optimization, efficiently solvable with modern solvers, such as Gurobi [15], to obtain global optimal solutions or determine infeasibility. This hierarchical structure results in a fast LGP framework, making it possible to be formulated in closed-loop fashion. The whole framework therefore exhibits reactive behaviors at both task and motion levels.
@@ III-A Target-centric LGP
The full path of LGP is composed of $K$ phases of a fixed duration $T$. Instead of checking feasibility and computing the cost of full path after finding the full skeleton candidate, we propose to use backward propagation from the target configuration. This requires a new target-centric LGP formulation:
@@
$h_{\text{switch}}(\bm{x}(kT)|a_{k},s_{k-1})=0$
$g_{\text{switch}}(\bm{x}(kT)\,|\,a_{k},s_{k-1})\leq 0$
@@
where $\bm{x}(t)$ and $\bm{u}(t)$ are the continuous trajectory and control variables, with $\bm{x}(kT)$ as the final configuration of phase $k$. $h_{\text{path}}$ and $g_{\text{path}}$ represent the path constraints, and $h_{\text{switch}}$, $g_{\text{switch}}$ denote the transition conditions between two phases with the action operator $a_{k}$. $g_{k}$ is the target to be achieved at phase $k$. $s_{k}\in\mathcal{G}(g_{k})$ and $h_{\text{goal}}(\cdot,\cdot)$ describes the goal specification logically and geometrically. $c(.)$ and $c_{T}(.)$ are the path cost and terminal cost to penalize control effort (e.g., energy, force) and the final geometric performance. The logical state transition from $s_{k-1}$ to $s_{k}$ is determined by $succ(.,.)$ as a function of $s_{k-1}$ and action $a_{k}\in\mathcal{A}(s_{k-1})$. $\mathcal{T}^{\dagger}(g_{k})$ identifies the essential subgoals to attain $g_{k}$, which is described in detail in Alg. 1. The formulation (1) adopts the concepts and variables from the standard LGP formulation [5], while allowing dynamic programming for backpropagation from the target configuration $g_{K}$. This mechanism can be viewed as spatial reasoning over the causal structure of the task.
@@ III-B Mixed-integer Motion Planning
The goal of tabletop rearrangement is to manipulate blocks to achieve a desired configuration. This task involves a set of objects denoted as $\mathcal{O}$, each with poses represented as $P_{n}^{t}=(x_{n}^{t},y_{n}^{t},\theta_{n}^{t})\in SE(2)$ at time $t$ for object $n$. In essence, the problem entails determining the optimal placement $\bm{u}_{t}=[x_{j}^{t}\;y_{j}^{t}]^{\scriptscriptstyle\top}$ for a selected object $j$. This task is notably challenging due to the non-convex nature of the solution set. Rather than attempting to solve the problem within this non-convex set, we adopt a strategy of decomposition into multiple convex sets. Subsequently, the constraints for each object $n\in 1,\dots,N$ can be succinctly expressed as combinations of four halfspaces (we assume the objects are square blocks), denoted by $i\in\{1,2,3,4\}$:
$C_{i}:$ $[a_{i}\quad b_{i}]\;\bm{R}(\theta_{n}^{t})^{\scriptscriptstyle\top}\;\bm{u}_{t}\geq c_{i},$ (2)
where $\bm{R}(\theta_{n}^{t})$ is a rotation matrix of the block frame and w.r.t. the global frame:
$\bm{R}(\theta_{n}^{t})=\begin{bmatrix}\text{cos}(\theta_{n}^{t})&-\text{sin}(\theta_{n}^{t})\\ \text{sin}(\theta_{n}^{t})&\text{cos}(\theta_{n}^{t})\end{bmatrix},$
and $a_{1}=1,b_{1}=0,c_{1}=l/2+x_{n,b}^{t}$, $a_{2}=-1,b_{2}=0,c_{2}=l/2-x_{n,b}^{t}$, $a_{3}=0,b_{3}=1,c_{3}=l/2+y_{n,b}^{t}$, $a_{4}=0,b_{4}=-1,c_{4}=l/2-y_{n,b}^{t}$, $[x_{n,b}^{t}\quad y_{n,b}^{t}]=[x_{n}^{t}\quad y_{n}^{t}]\bm{R}(\theta_{n}^{t})$. $l$ is the size of the square block.
As a consequence, the non-intersection constraints among these four halfspaces can be viewed as a disjunctive constraint $\lor_{i=1}^{4}C_{i}$, which can be further represented by combining constraints along with the introduction of additional binary variables, as discussed in [16].
The above formulation can be converted as Mixed-Integer Quadratic Program (MIQP) characterized by a quadratic objective and linear constraints. This transformation incorporates binary variables denoted as $z_{i}$ to determine the specific halfspace within which the placement should reside. In the case of each halfspace, the constraints can be expressed linearly. To articulate the if-else condition related to binary variables, we can employ a conventional big-M formulation, similar to [17], to represent (2) as
$C_{n}:$ $[a_{i}\quad b_{i}]\;\bm{R}(\theta_{n}^{t})^{\scriptscriptstyle\top}\;\bm{u}_{t}\geq c_{i}-M(1-z_{i}),$ (3)
$\sum z_{i}=1,\quad z_{i}\in\{0,1\},\quad\forall i\in\{1,2,3,4\}.$
Given the action $a_{k}$, the motion planning problem in (1) can be described as
@@
$\bm{u}(t)\in\cup_{n=1}^{N}\{C_{n}\}-\{C_{m}\},$
where $m$ represents the block intended to be positioned by action $a_{k}$. Additionally, another similar constraint defined by the geometric goal should be taken into account for the intermediate actions. For instance, as illustrated in Fig. 2b, the action Pick[B #P2] must ensure that block B does not occupy position #P3 to facilitate subsequent actions.